OpenStudy (anonymous):

Erica bought 3 1/2 yards of fabric. If she uses 2/3 of the fabric, how much would she have left?

OpenStudy (anonymous):

Total is : \(\huge 3\frac{1}{2}\)

OpenStudy (anonymous):

And you should know: \[\large a \frac{b}{c} = \frac{ac + b}{c}\]